OpenStudy (anonymous):

$HX\left(\frac{1}{\sqrt(2)}|0\rangle+\frac{1}{\sqrt(2)}|+\rangle\right)$

OpenStudy (anonymous):

\[\left(\frac{1}{\sqrt(2)}|0\rangle+\frac{1}{\sqrt(2)}|+\rangle\right)\]

OpenStudy (anonymous):

put \[ instead of $

OpenStudy (anonymous):

\[HX\left(\frac{1}{\sqrt(2)}|0\rangle+\frac{1}{\sqrt(2)}|+\rangle\right)\]

OpenStudy (anonymous):

is this quantum mechaanics

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you are not allowed to ask anything about your edX midterms online remember?

OpenStudy (anonymous):

I did not ask about the exam question

OpenStudy (anonymous):

okay but i am sry i cant help cos i am also taking the course but confused!!! i left qs cos i cud not mange with two courses just left with electricity

OpenStudy (anonymous):

how far are you

OpenStudy (anonymous):

so far I score 59%

OpenStudy (anonymous):

oh that very good i want to do it but it will disturb my college midterms so i will just do electricity

OpenStudy (anonymous):

I need to clarify some concept so I modify the question.

OpenStudy (anonymous):

y dont you check your notes or ask Turing Test

OpenStudy (anonymous):

@TuringTest